OpenStudy (anonymous):

What is the equation, in standard form, of a horizontal ellipse centered at (11, –6) with a major axis of length 16 and a minor axis of length 12?

OpenStudy (anonymous):

standard form for a horizontal ellipse is: \(\large \frac{(x-h)^2}{a^2}+\frac{(x-k)^2}{b^2}=1 \) where (h, k) is your center; 2a and 2b is the length of your major and minor axes, respectively. in your example, (h, k) is (11, -6); a = 8; b = 6 now just plug those in to get your equation....